OpenStudy (anonymous):

A movie rental store has a movie selection that consists of 4 action movies, 3 horror movies, 2 drama movies, and 3 comedies. If you select two movies at random, what is the probability that you select 2 action movies?

OpenStudy (anonymous):

take the number of action movies to the total number of movies and theres your ratio @ipringle

OpenStudy (anonymous):

k

OpenStudy (anonymous):

4/12

OpenStudy (kropot72):

Assuming the selection is made without replacement, This can be solved as follows: The probability of selecting an action movie on the first selection is 4/12. Having selected an action movie on the first selection, there are 3 action movies in the remaining 11 movies. Therefore the probability of selecting an action movie on the second selection is 3/11. The required probability is then given by the product of these two values of probability: \[\large P(2\ action\ movies)=\frac{4}{12}\times\frac{3}{11}=you\ can\ calculate\]